OpenStudy (anonymous):

I do not know how to approach these. I can go further in #2 but I have no idea how to even start #1. Both these limits are limx→Π/4 1. sinx−cosxtanx−1 2. (1tanx−1−2tan2x−1) And for this one I got 4 as the answer but the answer is 2 limx→4x−4x√−8−x−−−−−√

OpenStudy (anonymous):

I do not know how to approach these. I can go further in #2 but I have no idea how to even start #1. Both these limits are \[\lim_{x \rightarrow \frac{ \Pi }{ 4 } }\] 1. \[\frac{ \sin x-\cos x }{ \tan x-1 }\] 2. \[(\frac{ 1 }{ \tan x-1 }-\frac{ 2 }{\tan ^{2}x-1 })\] Another one which is different And for this one I got 4 as the answer but the answer is 2. 3. \[\lim_{x \rightarrow 4}\frac{ x-4 }{ \sqrt{x}-\sqrt{8-4} }\]

OpenStudy (ash2326):

We have \[\lim_{x\to \frac{\pi}{4}}\frac{ \sin x-\cos x }{ \tan x-1 }\] We know tan x=sin x \ cos x so \[\lim_{x\to \frac{\pi}{4}}\frac{ \sin x-\cos x }{ \sin x-\cos x } \times \cos x \] \[\lim_{x\to \frac{\pi}{4}}\frac{\cancel { \sin x-\cos x }}{\cancel{ \sin x-\cos x }} \times \cos x \] \[\lim_{x\to \frac{\pi}{4}} \cos x\] Can you find the limit now?

OpenStudy (ash2326):

@cali2 ???

OpenStudy (anonymous):

I think so....0.999?

OpenStudy (ash2326):

\[cos \frac{\pi} 4= \frac 1 {\sqrt 2}\]

OpenStudy (ash2326):

Do you get it @cali2 ?

OpenStudy (anonymous):

Oki, I got how you got the answer to cos pi/2. Can you break down your. I know tanx= sinx/cosx but I am not sure how you got to cancel

OpenStudy (ash2326):

numerator and denominator are same that's why I cancelled \[\lim_{x\to \frac{\pi}{4}}\frac{\cancel { \sin x-\cos x }}{\cancel{ \sin x-\cos x }} \times \cos x\]

OpenStudy (ash2326):

and it's cos pi/4 not pi/2

OpenStudy (anonymous):

That was a typo. Okay, Thank you

OpenStudy (ash2326):

Welcome, now you try the same things with other questions Try to cancel the common terms